the line integral    (1,2,3) (2,0,0) (xdx ydy zdz) is independent of path or not.

Grade:12th pass

3 Answers

Arun
4 years ago
Dear student
 
Question is not complete. Please check and repost the proper and complete question. You can also attach an image.
avinash kumar
4 years ago
1.
  1. Describe the Taylor’s theorem for functions of two variables Jacobeans.

  2. Evaluate ]dxdy )x y(dzdx )y z(dydz )x z([ S ∫∫ + + + + + where S: x2 +y2 +z2 = 1

  3. Find volume of the region bounded by the first octant section cut from the region inside the cylinder x2 +y2 =1 and by the planes y=0, z=0, x=y.

  4. In usual notations prove that, = ∫∫∫ R f dv 2 ∫∫ S dA.

  5. In usual notations prove that, = ∫∫∫ R f dv 2 ∫∫ S dA

  6. Evaluate 1 0 2 2 2 0 x y dxdy π .

  7.  

  8. Check whether the line integral + + (1,2,3) (2,0,0) (xdx ydy zdz) is independent of path or not. 

  9. Verify Stoke's theorem for v y i x j 3 3  and surface S: the circular disk = x 2 +y2 1, z=0.

Vikas TU
4 years ago
Dear student 
Question is not clear 
Please attach an image, 
We will happy to  help you 
Good Luck
Cheers